# GP-GS: Gaussian Processes for Enhanced Gaussian Splatting

### v3

@@ 4.2 Multi Output Gaussian Process
Problem Definition. We consider a MOGP regression problem, where the goal is to predict dense point clouds given image pixels and their corresponding depth priors. Given a set of RGB images =$\mathcal{I}=\{I_{i}\}_{i=1}^{n}$, where $n$ is the total number of images, we extract the corresponding sparse point clouds $\mathcal{P}=\{\mathbf{P}_{j}\}_{j=1}^{N}$, where $N$ is the number of 3D points, each point $\mathbf{P}_{j}$ in the point clouds contains both 3D spatial position and colour information: $\mathbf{P}_{j}=(x_{j},y_{j},z_{j},r_{j},g_{j},b_{j})$, where $(x_{j},y_{j},z_{j})$ are 3D coordinates and $(r_{j},g_{j},b_{j})$ are RGB grey values. However, one point in the 3D point clouds may correspond to multiple image pixels from various views. To deal with this, we build a correspondence between 2D pixels and 3D point clouds $f:\mathbf{V}_{i}\rightarrow\mathbf{P}_{j}$, where $\mathbf{V}_{i}=\left(u_{i},v_{i}\right)$, with $(u_{i},v_{i})$ the pixel coordinates in the image space. All pixel coordinates are denoted by $\mathcal{V}=\{\mathbf{V}_{i}\}_{i=1}^{\tilde{n}}$. When multiple pixels $\mathbf{V}_{i_{1}},\mathbf{V}_{i_{2}},...,\mathbf{V}_{i_{m}}$ contribute to the estimation of a single point $\mathbf{P}_{j}$, we denote the correspondence as
@@
With the observation that the depth of pixels contributes to the densification performance, we leverage depth priors $\mathcal{D}=\{D_{i}\}_{i=1}^{n}$ obtained from monocular depth estimation methods, such as (Yang et al. 2024b), to provide accurate estimated depth values $D_{i}(u_{i},v_{i})$ for each pixel with coordinate $(u_{i},v_{i})$. For brevity, we will use $D_{i}$ instead of $D_{i}(u_{i},v_{i})$ below. The rationale behind choosing the monocular depth estimation method is that it strikes a trade-off between accuracy and efficiency, which will accelerate the overall performance of the GP-GS framework.
MOGP Formulation. The input features $\mathbf{X}$ of the MOGP consist of pixel coordinates and depth information in each image defined as $\mathbf{x}_{i,j}=\big{(}u_{i},v_{i},D_{i}\big{)}$, where $(u_{i},v_{i})$ are pixel coordinates, and $D_{i}$ is the corresponding depth value. The output targets $\mathbf{Y}$ include the corresponding position coordinates and RGB values $\mathbf{y}_{j}=(x_{j},y_{j},z_{j},r_{j},g_{j},b_{j})$. Notably, $\mathbf{x}_{i,j}$ is the entry of $\mathbf{X}$ indexed by the subscripts $i$ and $j$, and $\mathbf{y}_{j}$ is the $j$-th entry of $\mathbf{Y}$.
We then define the MOGP model as:
$\mathbf{Y}\sim\mathcal{MOGP}\big{(}\mathbf{m}(\mathbf{x}),\mathbf{K}(\mathbf{x},\mathbf{x}^{\prime})\big{)},$ (8)
where $\mathbf{m}(\mathbf{x})$ is the mean function and $\mathbf{K}(\mathbf{x},\mathbf{x}^{\prime})$ is the covariance function:
$\mathbf{m}(\mathbf{x})=\begin{bmatrix}m_{1}(\mathbf{x})&m_{2}(\mathbf{x})&\dots&m_{6}(\mathbf{x})\end{bmatrix}^{T},$ (9)
$\mathbf{K}(\mathbf{x},\mathbf{x}^{\prime})=\begin{bmatrix}k_{11}(\mathbf{x},\mathbf{x}^{\prime})&\cdots&k_{16}(\mathbf{x},\mathbf{x}^{\prime})\\ \vdots&\ddots&\vdots\\ k_{61}(\mathbf{x},\mathbf{x}^{\prime})&\cdots&k_{66}(\mathbf{x},\mathbf{x}^{\prime})\end{bmatrix}.$ (10)
@@
Loss Function and Optimization. For MOGP training, we employ a combined loss function that integrates negative log marginal likelihood and L2 regularization as shown in equation (11). Such an optimisation approach balances data likelihood maximization and regularization, leading to a more robust and generalizable MOGP model.
$\mathcal{L}_{\text{total}}=-\log p(\mathbf{y}|\mathbf{X})+\lambda\sum_{i=1}^{N}\|\theta_{i}\|_{2}^{2},$ (11)
@@
where $p(\mathbf{y}|\mathbf{X})$ represents the marginal likelihood of the observed output given the inputs, $\lambda\sum_{i=1}^{N}\|\theta_{i}\|_{2}^{2}$ is the L2 regularization term, with $\lambda=10^{-6}$ representing the weight decay. The L2 regularization helps prevent overfitting by penalizing large weights in the model parameters $\theta$.
@@
where $\Gamma(\cdot)$ is the Gamma function, $\sigma^{2}$ is the variance, and $K_{\nu}(\cdot)$ is a modified Bessel function of the second kind. The parameter $\nu$ modulates how smooth the function can be: smaller $\nu$ values permit abrupt transitions, while larger $\nu$ values enforce smoother spatial changes. Details of tuning $\nu$ on various datasets can be found in Table 5. Our MOGP training procedure can be found in Appendix A.1 Algorithm 1.
4.3 Point Cloud Densification
The process of point cloud densification involves an adaptive neighbourhood-based sampling strategy, which selectively identifies pixels as candidate inputs for the MOGP model (in section 4.2). The MOGP model subsequently predicts additional point clouds and applies uncertainty filtering to enhance the density.
Adaptive Sampling Strategy. To adaptively generate sampling pixels within the image domain, we introduce an adaptive neighbourhood-based sampling mechanism that places samples in circular neighbourhoods surrounding each available training point $(u_{i},v_{i})$. Specifically, we define a set of $N$ sampled pixels $\mathcal{\tilde{P}}$ as follows:
$\quad\Bigm{|}\,(u_{i},\,v_{i})\in\mathcal{V}_{i}\Bigr{\}},$ (13)
where $\theta_{j}\in[0,2\pi)$ are uniformly distributed angles that control the sampling directions, $r=\beta\cdot\min(H,W)$ is the adaptive movement radius, $\beta\in(0,1)$ controls the sampling scale, $W$ and $H$ denote the image width and height, respectively. $W$ and $H$ are used to normalize all samples to the range $[0,1]$ in equation (13). The parameter $M$ determines the angular resolution of the sampling process. For each sampled pixel $(u_{i},v_{i})$, the corresponding depth value $D_{i}$ is retrieved from the depth image $\mathcal{D}$.
[…]
An overview of the uncertainty-based filtering procedure can be found in Appendix A.1 Algorithm 2.
5 Experiments
5.1 Dataset and Implementation Details
Dataset. Previous approaches to SfM often overlook the explicit correspondence between 2D image pixels and sparse 3D point clouds, we address this gap by introducing the Pixel-to-Point-Cloud dataset. This dataset provides a detailed mapping between image pixels and their corresponding 3D point clouds, thus facilitating more accurate modelling and enabling the training of our MOGP model. We further conduct a comprehensive evaluation of our method on dataset spanning NeRF Synthetic (Mildenhall et al. 2021), Mip-NeRF 360 (Barron et al. 2022), and Tanks and Temples (Knapitsch et al. 2017). These benchmarks are popular in 3D reconstruction and novel-view synthesis due to their diverse characteristics, including high-fidelity 3D models, large-scale outdoor scenes with intricate scenes like clustered leaves, and lighting condition variations. By leveraging our new dataset and evaluation framework, we demonstrate the effectiveness of our approach in advancing point cloud densification and improving reconstruction fidelity across varied and challenging scenarios.
Implementation Details The MOGP-based densification process in GP-GS can be regarded as a plug-and-play module and can be integrated into any 3D reconstruction framework based on SfM sparse point clouds. We have integrated the process into 3DGS*, one of the state-of-the-art 3D reconstruction models, and compared their performance over the datasets mentioned earlier following 3DGS*’s training hyperparameters. Our MOGP models are trained for 1,000 iterations across all scenes, we set L2 regularization weight $\lambda=10^{-6}$, learning rate $l=0.01$, dynamic sampling resolution $M=8$, adaptive sampling radius $r=0.25$. All experiments are conducted on an RTX 4080 GPU.
Metrics. The evaluation of the GP-GS is twofold, i.e., the evaluation of the SfM sparse point densification, and the evaluation of the novel view rendering performance. For densification, in line with prior studies on point cloud reconstruction and completion (Yu et al. 2021; Yuan et al. 2018), we adopt the mean Chamfer Distance (CD) to measure the discrepancy between the predicted point clouds and the ground truth. Specifically, for a predicted point set $\mathcal{P}$ and its corresponding ground truth set $\mathcal{G}$, the CD between them is calculated as:
We also compute the Root Mean Squared Error (RMSE) and $R^{2}$ (Edwards et al. 2008) to show the robustness of GP-GS under various metrics. The RMSE is calculated to quantify the average squared deviation between the predicted and true values, using $\mathrm{RMSE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(\mathbf{g_{i}}-\mathbf{p_{i}})^{2}}$, a lower RMSE value indicates a more accurate prediction. The $R^{2}$ score captures the proportionate reduction in the residual variance:
For novel view rendering performance evaluation, we compare our method against state-of-the-art NVS approaches based on commonly used image-based metrics: Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM) (Wang et al. 2004), and Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al. 2018) on the rendered images in the test views.
5.2 Results of Gaussian Processes
Figure 3: Qualitative comparison of 3D Gaussians during 3DGS training process. The left column shows the ground truth images. The middle column presents 3D Gaussians from 3DGS*, while the right column shows 3D Gaussians generated using our MOGP.
Quantitative Results To ensure the best performance of MOGP for densification, we evaluated MOGP model with different hyperparameters on our Pixel-to-Point-Cloud dataset, e.g., $\nu$ with values such as ($\nu\in\{0.5,1.5,2.5\}$). The data were split into an 80% training subset and a 20% testing subset, and the model was subsequently assessed using $R^{2}$, RMSE, and CD as we mentioned in section 5.1. The average outcomes of this analysis are summarized in Table 1, which shows that $\nu=0.5$ consistently yields the most favourable results. Consequently, we adopt $\nu=0.5$ as our default configuration for subsequent experiments. Detailed metrics across multiple datasets for this kernel selection procedure are presented in Appendix 5.2 Table 5.
Table 1: Comparison of different Matern kernel smoothness parameters $\nu$. The best results are highlighted.
$\nu$ Value $R^{2}\uparrow$ $\mathrm{RMSE}\downarrow$ $\mathrm{CD}\downarrow$
0.5 0.71 0.13 0.21
1.5 0.67 0.15 0.24
2.5 0.61 0.14 0.26
To validate the effectiveness of our variance-based filtering algorithm, we analyze the variance distributions across multiple Pixel-to-Point datasets, including NeRF Synthetic, Tanks & Temples, and Mip-NeRF 360, due to their diverse characteristics. The reduction in variance in all datasets suggests that our filtering method generalizes well to different types of 3D scenes, improving overall point cloud quality and enabling more accurate scene representation, as shown in Figure 5, the left column shows the spatial and colour variances before filtering on Tanks & Temples (Truck), the right column presents the variances after applying our filter. The variances in spatial components $(x,y,z)$ and colour channels $(r,g,b)$ are significantly reduced, demonstrating that our filtering process effectively removes high-variance noisy points while preserving scene consistency. More results on various datasets can be found in Appendix A.3.
Figure 4: Qualitative results of our point cloud densification approach.
Figure 5: Variance reduction comparison of our variance-based filtering algorithm on Tanks & Temples (Truck).
Further validating our densified point cloud, we present Figure 3, which shows 3D Gaussians during training. Our approach leads to more structured and coherent reconstructions, especially in complex regions with fine textures and occlusions. In the flower scene, our 3D Gaussian representation successfully reconstructs branches, trunks, and leaves on the left and right sides, while the original 3D Gaussian sphere remains blurry and lacks these details. Additionally, our method accurately preserves the flower shape in the middle, improving overall reconstruction fidelity. In the room scene, our method provides a clearer representation of the curtain folds on the left, whereas the original method produces irregularly shaped Gaussian spheres and visible artefacts. The highlighted insets further demonstrate that our method reduces visual artefacts and better preserves geometric consistency during training.
5.3 Quantitative and Qualitative Results on NVS
Figure 6: Qualitative comparison of ours GP-GS and 3DGS*
To evaluate our method’s effectiveness in novel view synthesis, we conduct comprehensive experiments on datasets like Tanks & Temples, Mip-NeRF 360, and NeRF Synthetic. We have compared the proposed GP-GS with the state-of-the-art 3DGS* using both quantitative metrics and qualitative visualizations.
Quantitative Results Table 2,  3,  4 report PSNR, SSIM, and LPIPS scores across all datasets, respectively. Our method consistently achieves higher PSNR and SSIM values while maintaining lower LPIPS scores, indicating superior reconstruction quality and perceptual realism. The improvements are particularly evident in regions with high structural complexity and varying textures, such as dense foliage and intricate surfaces, where existing methods often suffer from over-smoothing and aliasing artefacts. Our approach effectively preserves fine details and enhances geometric consistency, even under challenging lighting conditions. Figure 6 provides a detailed visual comparison.
Table 2: Comparison of Various Methods on Tanks and Temples
Dataset 3DGS* Ours
[…]
Truck 25.10 0.876 0.153 25.18 0.877 0.153
Train 22.13 0.819 0.202 22.22 0.823 0.199
Average 23.62 0.847 0.178 23.7 0.85 0.176
Table 3: Comparison of Various Methods on NeRF Synthetic
Dataset 3DGS* Ours
[…]
chair 30.99 0.978 0.024 30.99 0.978 0.024
drums 25.52 0.940 0.065 25.52 0.940 0.065
ficus 25.03 0.923 0.058 25.20 0.923 0.058
hotdog 34.53 0.979 0.037 34.55 0.979 0.037
lego 28.32 0.929 0.074 29.08 0.936 0.067
materials 16.82 0.734 0.239 16.68 0.736 0.241
mic 18.90 0.855 0.144 19.02 0.863 0.123
ship 27.08 0.848 0.157 27.05 0.849 0.154
Average 25.90 0.898 0.100 26.01 0.901 0.096
Table 4: Comparison of Various Methods on Mip-NeRF 360
Dataset 3DGS* Ours
[…]
bicycle 19.169 0.632 0.248 19.219 0.634 0.248
bonsai 18.66 0.579 0.458 18.69 0.581 0.455
counter 29.15 0.911 0.122 29.17 0.915 0.121
garden 29.10 0.915 0.063 29.38 0.920 0.061
kitchen 31.74 0.943 0.069 32.28 0.948 0.068
room 31.78 0.934 0.126 31.89 0.939 0.121
flowers 23.04 0.707 0.251 23.16 0.711 0.246
Average 26.09 0.803 0.191 26.26 0.807 0.189
Qualitative Results As shown in Figure 6, we compare our method with 3DGS* on datasets Tanks & Temples (Truck), Mip-NeRF 360 (Flowers and Bicycles), and NeRF Synthetic (Lego), we provide additional results in Appendix A.4 for each scene to further substantiate our conclusions. Our method demonstrates superior performance in NVS by effectively reconstructing fine-grained details, reducing artifacts, and enhancing occlusion handling. The visual comparisons highlight key challenges in 3DGS* rendering, including occlusions, high-frequency textures (e.g., object edges and surface details), and lighting variations. Notably, the zoom-in regions (marked with red and blue boxes) illustrate our model’s ability to recover fine structures more accurately than existing approaches. These results further validate the robustness of our method in handling complex scenes, offering improved perceptual quality and structural consistency across diverse datasets.
6 Conclusion
We propose GP-GS, a novel framework that enhances 3DGS initialization and improves rendering quality by addressing the limitations of sparse SfM reconstructions. Our method formulates point cloud densification as a continuous regression problem using MOGP, which adaptively densifies sparse SfM point clouds by learning mappings from 2D image pixels and depth priors to 3D positions and colours. We introduce an adaptive neighbourhood-based sampling strategy for structured densification and apply variance-based filtering to remove high-uncertainty predictions. Experiments on synthetic and real-world datasets show that GP-GS significantly improves 3D reconstruction quality, especially in densely packed or challenging lighting conditions. Future work includes integrating temporal information for dynamic scenes and scaling to larger environments.
Appendix A Appendix
The Appendix provides additional details on our MOGP model training and testing. It also includes additional experimental results that are omitted from the main paper due to space constraints.
A.1 MOGP Training and Dynamic Filtering Procedure
This section provides details on the training, dynamic sampling and uncertainty-based filtering process of our MOGP model for point cloud densification. The full MOGP training procedure is summarized in Algorithm 1, and the dynamic sampling and filtering process is detailed in Algorithm 2.
@@ A.1 MOGP Training and Dynamic Filtering Procedure
$\mathcal{L}_{\text{total}}=-\log p(\mathbf{Y}|\mathbf{X})+\lambda\sum_{i=1}^{N}\|\theta_{i}\|_{2}^{2}$ (18)
@@
  • Initialize: $\theta=\{\sigma,\nu,\dots\}$.
@@
    • Compute gradient $\nabla_{\theta}\mathcal{L}_{\text{total}}$.
@@
      $\theta\leftarrow\theta-\eta\,\nabla_{\theta}\,\mathcal{L}_{\text{total}}$ (19)

### v4

@@ IV-B Multi Output Gaussian Process
Problem Definition. We consider a MOGP regression problem, where the goal is to predict dense point clouds given image pixels and their corresponding depth priors. Given a set of $n$ RGB images =$\mathcal{I}=\{I_{i}\}_{i=1}^{n}$, we can extract the corresponding sparse point clouds $\mathcal{P}=\{\mathbf{P}_{j}\}_{j=1}^{N}$ with a total number of $N$ points. Each point $\mathbf{P}_{j}$ in the point clouds contains its 3D spatial position and the associated color information: $\mathbf{P}_{j}=[x_{j},y_{j},z_{j},r_{j},g_{j},b_{j}]^{T}$, where $[x_{j},y_{j},z_{j}]^{T}$ are 3D coordinates and $[r_{j},g_{j},b_{j}]^{T}$ are RGB values.
Since a single 3D point can be visible from multiple views, it is straightforward to say that there can be multiple pixels from different images correspond to the same 3D point. Therefore, it is critical to learn the correspondences between 2D pixels and 3D points. To achieve this, we first perform pixel-to-point cloud matching to establish the correspondence between pixels and 3D points. This is done by projecting the sparse point clouds back to the image plane of each view and finding the corresponding pixel coordinates. The matching process can be formulated as:
@@
where $\mathbf{V}_{i_{k}}$ are the pixels across different views $k=1,\cdots,m$ that project to the same point $\mathbf{P}_{j}$. Using $f$, we evaluate the contribution of RGB images =$\mathcal{I}=\{I_{i}\}_{i=1}^{n}$ to point clouds $\mathcal{P}=\{\mathbf{P}_{j}\}_{j=1}^{N}$ based on the number pixel-to-point correspondences. The RGB image holds the most correspondences will be selected as the key frame for MOGP training.
@@
To enhance densification performance, we leverage the key frame depth priors obtained from monocular depth estimation methods, such as [36]. These depth priors provide estimated depth values $\mathcal{D}_{i}(u_{i},v_{i})$ for each pixel with coordinate $[u_{i},v_{i}]^{T}$. For brevity, we will use $\mathcal{D}_{i}$ instead of $\mathcal{D}_{i}(u_{i},v_{i})$ below. The rationale behind choosing the monocular depth estimation method is that it strikes a trade-off between accuracy and efficiency, which will accelerate the overall performance of the GP-GS framework.
MOGP Formulation. Let the input features of the MOGP be denoted as $\mathbf{X}_{i}=[u_{i},v_{i},\mathcal{D}_{i}]^{T}\in\mathbb{R}^{3}$, where $(u_{i},v_{i})$ are the pixel coordinates and $\mathcal{D}_{i}$ is the associated depth value from a keyframe image. The corresponding output targets are $\mathbf{Y}_{i}=[x_{i},y_{i},z_{i},r_{i},g_{i},b_{i}]^{T}\in\mathbb{R}^{6}$, representing the 3D spatial coordinates and RGB values. For notational convenience, we use $\mathbf{x}:=\mathbf{X}_{i}\in\mathbb{R}^{3}$ and $\mathbf{y}:=\mathbf{Y}_{i}\in\mathbb{R}^{6}$ to denote a single input-output pair.
We define a Multi-Output Gaussian Process (MOGP) over the function $\mathbf{y}(\cdot):\mathbb{R}^{3}\rightarrow\mathbb{R}^{6}$ as:
$\mathbf{y}(\mathbf{x})\sim\mathcal{GP}\left(\mathbf{m}(\mathbf{x}),\mathbf{K}(\mathbf{x},\mathbf{x}^{\prime})\right),$ (9)
where the mean function $\mathbf{m}(\mathbf{x})\in\mathbb{R}^{6}$ is defined as:
$\mathbf{m}(\mathbf{x})=\begin{bmatrix}m_{1}(\mathbf{x})\\ m_{2}(\mathbf{x})\\ \vdots\\ m_{6}(\mathbf{x})\end{bmatrix},\quad\text{with }m_{j}:\mathbb{R}^{3}\rightarrow\mathbb{R},$ (10)
and the covariance function $\mathbf{K}(\mathbf{x},\mathbf{x}^{\prime})\in\mathbb{R}^{6\times 6}$ is given by:
$\mathbf{K}(\mathbf{x},\mathbf{x}^{\prime})=\begin{bmatrix}k_{11}(\mathbf{x},\mathbf{x}^{\prime})&\cdots&k_{16}(\mathbf{x},\mathbf{x}^{\prime})\\ \vdots&\ddots&\vdots\\ k_{61}(\mathbf{x},\mathbf{x}^{\prime})&\cdots&k_{66}(\mathbf{x},\mathbf{x}^{\prime})\end{bmatrix},$ (11)
with $k_{ij}:\mathbb{R}^{3}\rightarrow\mathbb{R}$.
@@
Loss Function and Optimization. For MOGP training, we employ a combined loss function that integrates negative log marginal likelihood and L2 regularization as shown in Equation (12). Such an optimization approach balances data likelihood maximization and regularization, leading to a more robust and generalizable MOGP model.
$\mathcal{L}_{\text{total}}=-\log p(\mathbf{Y}|\mathbf{X})+\lambda\|\boldsymbol{\theta}\|_{2}^{2},$ (12)
@@
where $p(\mathbf{Y}|\mathbf{X})$ represents the marginal likelihood of the observed output given the inputs, $\lambda\|\boldsymbol{\theta}\|_{2}^{2}$ is the L2 regularization term, with $\lambda=10^{-6}$ representing the weight decay. The L2 regularization helps prevent overfitting by penalizing large model parameters $\theta_{i}\in\boldsymbol{\theta}$, which are determined by the kernel.
@@
where $\Gamma(\cdot)$ is the Gamma function, $\sigma^{2}$ is the variance, and $K_{\nu}(\cdot)$ is a modified Bessel function of the second kind. The parameter $\nu$ modulates how smooth the function can be: smaller $\nu$ values permit abrupt transitions, while larger $\nu$ values enforce smoother spatial changes. Details of tuning $\nu$ on various datasets can be found in Table II. Our MOGP training procedure can be found in Algorithm 1.
@@
$\mathcal{L}_{\text{total}}=-\log p(\mathbf{Y}|\mathbf{X})+\lambda\|\boldsymbol{\theta}\|_{2}^{2}$ (16)
@@
  • Initialize: $\boldsymbol{\theta}=[\sigma,\nu,\dots]^{T}$.
@@
    • Compute gradient $\nabla_{\boldsymbol{\theta}}\mathcal{L}_{\text{total}}$.
@@
      $\boldsymbol{\theta}\leftarrow\boldsymbol{\theta}-\eta\,\nabla_{\boldsymbol{\theta}}\,\mathcal{L}_{\text{total}}$ (17)
@@
IV-C Point Cloud Densification
Point cloud densification involves an adaptive neighbourhood-based sampling strategy, which selectively identifies pixels as candidate inputs for the MOGP model (in section IV-B). The MOGP model subsequently predicts additional point clouds and applies uncertainty filtering to enhance the density.
Adaptive Sampling Strategy. To adaptively generate sampling pixels within the image domain, we introduce an adaptive neighbourhood-based sampling mechanism that samples from immediate circular neighbourhoods of each available training point $(u_{i},v_{i})$. Specifically, we define a set of $N$ sampled pixels $\mathcal{\tilde{P}}$ as follows:
$\quad\Bigm{|}\,(u_{i},\,v_{i})\in\mathcal{V}_{i}\Bigr{\}},$ (18)
where $\theta_{j}\in[0,2\pi)$ are uniformly distributed angles that control the sampling directions, $r=\beta\cdot\min(H,W)$ is the adaptive movement radius, $\beta\in(0,1)$ controls the sampling scale, $W$ and $H$ denote the image width and height, respectively. $W$ and $H$ are used to normalize all samples to the range $[0,1]$ in Equation (18). The parameter $M$ determines the angular resolution of the sampling process. For each sampled pixel $(u_{i},v_{i})$, the corresponding depth value $\mathcal{D}_{i}$ is retrieved from the depth image $\mathcal{D}$.
[…]
An overview of the uncertainty-based filtering procedure can be found in Algorithm 2.